OpenStudy (anonymous):

What is the proof of number 5? Given: 4x + 1 = 6x - 2, Prove: x = 3/2 - (1) 4x + 1 = 6x - 2 (2) 4x + 3 = 6x (3) 3 = 2x (4) 3/2 = x (5) x = 3/2

OpenStudy (jiteshmeghwal9):

put the value of x in the given equation & then try to solve u will gt L.H.S.=R.H.S

OpenStudy (jiteshmeghwal9):

& it will be the proof that x=3/2

OpenStudy (jiteshmeghwal9):

can u do it @jorgesnz ???????????

OpenStudy (anonymous):

not really, i don't get it :S

OpenStudy (asnaseer):

do you mean how do you get from step 4 to step 5?

OpenStudy (jiteshmeghwal9):

\[4 \times \frac{3}{2}+1=6 \times \frac{3}{2}-2\]

OpenStudy (jiteshmeghwal9):

\[2 \times 3+1=3\times 3-2 \]

OpenStudy (jiteshmeghwal9):

6+1=9-2 7=7 L.H.S=R.H.S

OpenStudy (jiteshmeghwal9):

hence it is proved that the value of x=3/2 is right

OpenStudy (jiteshmeghwal9):

do u gt it now @jorgesnz ?????

OpenStudy (anonymous):

@asnaseer No, is to find the reason of number 5

OpenStudy (anonymous):

@jiteshmeghwal9 , yeah

OpenStudy (jiteshmeghwal9):

nice to hear :)

OpenStudy (asnaseer):

do you understand each step from 1 to 5?